OpenStudy (abbie):

Kim walked to the grocery store and back home.The grocery store is 2 5/8 miles from her house. The next day kim walked twice as far to the post office and then back home.How far did she walk in all?

OpenStudy (agent0smith):

\[\large 2 \frac{ 5 }{ 8 }+ 2*(2 \frac{ 5 }{ 8 })\]

OpenStudy (abbie):

\[A. 7 \frac{ 7 }{ 8 } B. 5 \frac{ 1 }{ 4 } C. 10 \frac{ 1 }{ 2 } D. 15 \frac{ 3 }{ 4 }\]

OpenStudy (abbie):

2 5/8 + 4 5/8?

OpenStudy (agent0smith):

you have to multiply the 5 by 2 also

OpenStudy (abbie):

4 10/8

OpenStudy (abbie):

6 3/2 ?

OpenStudy (agent0smith):

\[\large 2 \frac{ 5 }{ 8 }+ 2*(2 \frac{ 5 }{ 8 }) = 2 \frac{ 5 }{ 8 } + 4 \frac{ 10 }{ 8 }\]\[\large = 2+ 4 + \frac{ 5 }{ 8 } + \frac{ 10 }{8 }\]

OpenStudy (agent0smith):

\[\large = 6 \frac{ 15 }{ 8 } = 7 \frac{ 7 }{8 }\]

OpenStudy (abbie):

thats what I put thnx for helping!

OpenStudy (agent0smith):

Good job :)

OpenStudy (abbie):

I thought I was wrong